# Personalized Generative Low-light Image Denoising and Enhancement

### v2

@@ 3.2 Why Combining Gallery and Physical Buffer?
  • Exp 1: Model trained solely on generic images (purely generative).
@@
  • Exp 2: Model trained on generic images, conditioned on the facial physical buffer extracted from the input image.
@@
  • Exp 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
@@
  • Exp 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer.
@@
We tested these models on a mildly noisy input. Exp 1 produces random faces, as expected from a purely generative model. In contrast, Exp 2 uses physical buffers obtained via DECA [15], which can extract coarse face buffers from a single 2D photo (see Fig. 4), yielding faces with user-like features (e.g., nose shape, mouth and skin tone). Hence, physical buffers can convey key identity cues. Exp 3 outputs distinct faces of the same user, compared to Exp1, showing that fine-tuning with gallery photos effectively transfers identity information. Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
@@
While Fig. 3 shows us a design direction, we need to overcome:
@@
    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user. Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
@@
    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [15]. Severe degradation undermines the extraction of physical buffers (Fig. 4), resulting in poor reconstruction. Moreover, features captured from a single degraded photo are limited, not to mention from a degraded one.
In conclusion, to avoid repeated fine-tuning on galleries and reduce the risk of erroneous buffers from noisy inputs, it is more desirable to jointly employing gallery photos and facial physical priors, as we will discuss below.
3.3 Extracting Physical Buffer from Photo Gallery
Based on the preceding analysis, we propose extracting physical buffers directly from existing clean gallery photos and using them as conditions to guide the diffusion model during training. This approach offers significant benefits: clean gallery photos allow us to form comprehensive and robust identity-consistent physical buffers containing more acurate facial identity features of the target person. Specifically, We will extract the albedo and normal maps of the target person, which are two key attributes used to describe the physical characteristics of a face.
  • Albedo captures the intrinsic color of an object’s surface—that is, the subject’s color properties without any lighting effects. It reveals a person’s facial appearance, e.g. eyebrows, lip coloration, and other distinct texture elements can reflect individual identity traits.
  • Normal encodes the surface orientation at each pixel, which provides hints about the underlying shape characteristics.
@@ 3.3 Extracting Physical Buffer from Photo Gallery
By conditioning on these identity physical buffers, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
@@
However, extracting physical buffers for a person from a set of gallery photos is a more challenging task than extracting these buffers from a single input, as in methods like DECA. This complexity arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency. As shown in the physical buffer extractor module in Fig. 5(a), we adopt the aggregation net design from the LAP [65] to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
3.4 Model Framework
We name our final proposed model DiffPGD. As illustrated in Fig. 5, the model leverages ID physical buffers (albedo $a^{i}$ and normal $n^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training directs the generative process toward a personalized space.
The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation. To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed. Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i, each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$ and $\mathbf{\alpha}_{k}^{n}$, respectively. To derive a global representation of the identity for person i based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted, This method learns a weight set $\{\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}\}_{k=1}^{N}$ according to the quality of each image in$\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$. The combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{w}_{k}^{a}\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{w}_{k}^{n}\mathbf{\alpha}_{k}^{n}.$ (3)
@@ 3.4 Model Framework
Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$.
During inference, the process begins with a Gaussian noise image $\mathbf{x}_{T}^{i}$, which is iteratively transformed into the clean, enhanced image $\mathbf{x}_{0}^{i}$ by sampling from $p(\mathbf{x}_{0}^{i}|\mathbf{x}_{T}^{i},\mathbf{y}^{i},\mathbf{a}^{i},\mathbf{n}^{i})$ using the trained model. With the low-light noisy image $\mathbf{y}^{i}$ as another condition during training, other features, such as pose, hairstyle, and background inside the image beyond facial features can be effectively captured as well.
@@
4 Experiments
Figure 6: Visual comparison on real cases. The first column presents the input images along with their capture parameters. In the second column, we enhance the brightness to better reveal the noise. Please zoom in for the best visual experience.
4.1 Dataset
Training Data. Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [30], the InverseISP algorithm [34], and the Poisson-Gaussian noise model Eq. 1. The pipeline involves converting RGB images to their corresponding RAW mosaiced format using InverseISP [34], followed by applying Eq. 1 to simulate low-light noise using realistic camera sensor parameters. Multiple noise levels are generated by varying the light intensity (Lux) to ensure the model is robust across diverse noise conditions. Refer to supplementary for more details on the simulation.
  Photons Per Pixel 5 10 13 26
Methods ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$ ID Score $\uparrow$ FID $\downarrow$ KID $\downarrow$ PSNR $\uparrow$
SNR-aware[59] 0.2156 188.3311 0.1302 8.4541 0.4240 139.8193 0.0786 9.5462 0.4999 124.1162 0.0626 10.1712 0.6937 99.9875 0.0383 12.4961
MIRNet[61, 62] 0.1229 294.7381 0.2414 8.6008 0.2860 214.7785 0.1534 9.8572 0.3830 187.2571 0.1286 10.4209 0.6212 136.1320 0.0793 12.8992
DiffLL[22] 0.2910 167.3905 0.1068 17.3506 0.4544 150.4559 0.1117 18.3360 0.5095 140.2738 0.0996 19.9547 0.6833 119.2693 0.0760 23.3180
GFPGAN[52] 0.1941 388.9293 0.4256 7.1990 0.2075 233.3173 0.1521 8.1283 0.2976 190.7264 0.1326 9.5483 0.5976 158.9475 0.0829 11.4855
CodeFormer[67] 0.2863 111.4544 0.04908 18.2406 0.4940 91.1244 0.0326 19.2692 0.5563 84.7955 0.0281 20.2306 0.7338 80.6582 0.0263 24.5512
Ours 0.3714 77.5743 0.0084 17.9334 0.5225 71.8645 0.0063 18.4999 0.5938 69.5567 0.0042 19.0910 0.7118 59.1975 0.0030 23.7226
Table 1: Quantitative comparison with state-of-the-art works on CelebAMask-HQ dataset on 5PPP, 10PPP, 13PPP, and 26PPP. The lower the PPP, severe the degradation is.
Testing Data. For testing DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
The simulation setup assumes a $12$-bit image sensor with parameters similar to that of Canon DSLR cameras: a Full Well Capacity of $36000\text{e}^{-}$, Quantum Efficiency of $0.42$, dark current of $7\text{e}^{-}$, and read noise of $6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance.

### v3

@@ III-B Why Combining Gallery and Physical Buffer?
  • Experiment 1: Model trained solely on generic images from the FFHQ dataset [Karras2018ASG].
@@
  • Experiment 2: Model trained on generic images, while conditioned on the facial physical buffer extracted from the noisy input image.
@@
  • Experiment 3: Model trained on generic images, then fine-tuned with the user’s gallery photos.
@@
  • Experiment 4: Model trained on generic images, fine-tuned with the user’s gallery photos, and conditioned on the facial physical buffer extracted from the input image.
@@
We tested these trained models on a mildly low-light noisy photo of a user, and the results are presented in the Fig. 3 (right). The model from Exp 1 produces random faces, as expected from a purely generative purpose. In contrast, the model from Exp 2 uses some coarse face physical buffers extracted from the input image [feng2021learning] as condition. The result from Exp 2 yields faces with user-like features (e.g., nose shape, mouth shape and skin tone). Hence, even just some coarse physical buffers extracted from a degraded photo can convey some key identity cues. Exp 3 outputs distinct faces of the same user, compared to Exp 1, showing that fine-tuning with gallery photos transfers identity information. Finally, Exp 4 yields the best results, showcasing that the collaborative use of gallery photos and facial physical prior can achieve the optimal outcomes.
Figure 3: Incorporating gallery photos and physical buffers enhances identity preservation. Exp 1: Trained on generic images only $\rightarrow$ outputs a generic face. Exp 2: Adds physical buffer conditioning $\rightarrow$ captures some user-specific features. Exp 3: Fine-tuned with gallery photos $\rightarrow$ generates an arbitrary user face. Exp 4: Combines gallery fine-tuning and physical buffer $\rightarrow$ restores the expected user face.
III-C Extracting Physical Buffer from Photo Gallery
@@
While Fig. 3 shows us a design direction, there are still two fundamental limitations we need to overcome:
@@
    Fine-tuning: Extracting identity via fine-tuning requires significant time and must be repeated for each user [ding2024restoration]. Additionally, the small size of a gallery compared with a large pre-trained model can lead to overfitting.
@@
    Unreliable buffers from degraded inputs: As studied in the face reconstruction literature, low-quality inputs consistently present additional challenges [feng2021learning]. Fig. 4 shows that extracting albedo and normal maps directly from degraded inputs becomes unreliable as the noise level increases, often leading to incorrect physical buffers, resulting in poor reconstruction. Moreover, features captured from a single image are coarse and limited, not to mention from a degraded one.
Our approach: Extract Physical Buffer from the gallery. Based on the preceding analysis, we propose extracting physical buffers directly from the clean gallery photos instead of noisy input image. This approach offers significant benefits: The gallery provides diverse but unstructured identity evidence, while the physical buffer offers structured, lighting-independent representations. By deriving the physical buffers from the clean gallery, we combine these complementary strengths—retaining the gallery’s diversity while achieving physically consistent and accurate facial identity features of the target person. The resulting ID-consistent buffers remain reliable even under severe noise and serve as a strong condition for the diffusion-based restoration process.
@@
By conditioning on the identity physical buffers directly from existing clean gallery photos, we eliminate the need to fine-tune the model with gallery images and avoid the issue of incorrect buffer extraction. Consequently, the model is better equipped to preserve the target person’s identity and facial features, even under severe degradation.
Figure 4: For input images affected by mild low-light noise, physical buffers can be accurately extracted from input. However, when encountering the input images degraded by severe low light and noise, precise physical buffers are difficult to obtain from input.
@@
How to extract physical buffer from gallery? However, extracting physical buffers from a set of gallery photos is more challenging than extracting from a single input. The complication arises because gallery photos of the same person are often taken under varying conditions, such as different scenes and poses. To ensure these buffers comprehensively represent the person’s identity and facial properties, they must be aligned to maintain identity consistency. Inspired by LAP [Zhang2021lap], we used an aggregation network as the physical buffer extractor module (shown in Fig. 5, to extract the ID-consistent physical buffers, which adaptively combines facial features from multiple images to learn consistent geometry and texture representations, and generates ID-consistent albedo and normal maps from a photo collection of the same individual.
Details of the proposed scheme. As illustrated in Fig. 5, the physical buffer extractor gets the ID physical buffers (albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$) extracted from person i’s gallery photos. The aggregation network contains a shared encoder $\delta$ across multiple images and a global decoder $\phi$ for predicting the consistent representation. To model albedo and normal, two separate aggregation networks, denoted as ($\delta^{a},\phi^{a}$) and ($\delta^{n},\phi^{n}$), are employed (Fig. 5 uses the albedo extraction as example, and the normal extraction follows the similar strategy).
Given a photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i (we will omit the identity index i in the following for simplicity), each image is processed by $\delta^{a}$ and $\delta^{n}$ to extract its texture and geometry latent codes $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$ $\in$ $[1,1,c]$, respectively. Due to the quality of $\{\textbf{F}_{k}\}_{k=1}^{N}$, the importance of each dimension in $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$ which reveals how correlated it is to the identity, is supposed to be different. Therefore, to derive a global representation of the identity based on $\{\mathbf{\alpha}_{k}^{a},\mathbf{\alpha}_{k}^{n}\}_{k=1}^{N}$, an adaptive aggregation method is adopted. More specifically, channel-wise weights $\mathbf{w}_{k}^{a},\mathbf{w}_{k}^{n}$ $\in$ $[1,1,c]$ are also learned to represent the importance of each dimension in $\mathbf{\alpha}_{k}^{a}$, $\mathbf{\alpha}_{k}^{n}$:
[…]
A soft-max function is further used to normalize the weights into $\{\hat{\mathbf{w}}_{k}^{a}\}_{k=1}^{N},\{\hat{\mathbf{w}}_{k}^{n}\}_{k=1}^{N}$. Finally, the combined global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ for texture and depth are calculated as:
$\mathbf{\alpha}_{c}^{a}=\sum_{k=1}^{N}\mathbf{\hat{w}}_{k}^{a}\cdot\mathbf{\alpha}_{k}^{a},\text{ and}\quad\mathbf{\alpha}_{c}^{n}=\sum_{k=1}^{N}\mathbf{\hat{w}}_{k}^{n}\cdot\mathbf{\alpha}_{k}^{n}.$ (8)
@@
Once calculated, the global ID-code $\mathbf{\alpha}_{c}^{a}$ and $\mathbf{\alpha}_{c}^{n}$ are passed to the decoders $\phi^{a}$ and $\phi^{n}$, respectively, to produce the ID-consistent albedo $\mathbf{a}$ and normal $\mathbf{n}$:
$\mathbf{a}=\phi^{a}(\mathbf{\alpha}_{c}^{a}),\text{ and}\quad\mathbf{n}=\phi^{n}(\mathbf{\alpha}_{c}^{n}).$ (9)
The algorithm for our physical buffers extraction process is shown in Algorithm 1.
Figure 5: ID-consistent physical buffer extractor. The extractor gets the ID physical buffers extracted from target person’s gallery photos. It contains a shared encoder across multiple images and a global decoder for predicting the consistent representation. Note we only show albedo here, and the normal extractor follows the similar strategy. The module weights are initialized from the aggregation network in [Zhang2021lap] and frozen during training.
Algorithm 1 Extracting physical buffers from photo gallery
1:Input: A photo gallery of N images $\{\textbf{F}_{k}^{i}\}_{k=1}^{N}$ of person i (we omit the identity index i in the following for simplicity),
[…]
6:return $\mathbf{a},\mathbf{n}$
Figure 6: Overview of our method. We leverage ID-consistent physical buffers (albedo and normal) extracted from gallery photos to constrain the generative space during diffusion restoration. The albedo encodes intrinsic skin/appearance while the normal represents facial geometry. In our framework, the output physical buffers isolate the intrinsic ID properties from lighting, shading, and pose, enabling the diffusion model to apply only ID-related information consistently.
III-D Diffusion-based Restoration Model Framework
In Section III-C we discussed how to extract the physical buffer. Now, we discuss how this physical buffer can be used in the restoration pipeline. We name our final proposed model DiffPGD, which adopts a diffusion-based restoration framework. As illustrated in Fig. 6, the model leverages ID physical buffers (albedo $\mathbf{a}^{i}$ and normal $\mathbf{n}^{i}$) extracted from person i’s gallery photos as conditions to guide our diffusion model during training to direct the generative process toward a personalized space. We will drop the superscript $i$ in the following for simplicity.
In restoration tasks, such as low-light denoising task solved in this paper, the goal is to recover a high-quality image $\mathbf{x}_{0}$ from a low-quality input $\mathbf{y}$. Unlike a pure generation task, where the output can be any image that follows the distribution of clean images, restoration is constrained by the content of $\mathbf{y}$. As shown in the conditional denoising diffusion model in Fig. 6, the forward/diffusion process (from right to left) gradually adds Gaussian noise, denoted by $q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$. Our goal is to reverse the diffusion process (from left to right) by gradually recovering the image from the input Gaussian noise with conditions from different modalities, which corresponds to learning the reverse process of a fixed Markov Chain of length $T$ conditioned on $\mathbf{y}$ and the facial physical buffers albedo and normal $(\mathbf{a},\mathbf{n})$, extracted from the photo gallery. When training our model, we adopt a P2-weighted loss format [choi2022perception] with an additional weight term $\lambda_{t}$, and the corresponding objective function measures the discrepancy between the predicted and ground-truth noise, and can be formatted as:
$\mathcal{L}=\lambda_{t}\mathbb{E}_{\mathbf{x}_{0},\mathbf{y},\mathbf{a},\mathbf{n},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n})\|_{2}^{2}],$ (10)
According to Eq. (10), $\epsilon_{\theta}$ takes as input the noisy target image $\mathbf{x}_{t}$, the time step $t$, and the conditional signal $\mathbf{c}$, which consists of the low-quality input image $\mathbf{y}$ and the extracted physical buffers $(\mathbf{a},\mathbf{n})$, to estimate the noise $\epsilon$. $\lambda_{t}$ serves as a hyperparameter to adjust the loss weight dynamically across different time step. It assigns higher weights to the loss at levels where the model learns perceptually rich contents while minimal weights to which the model learns imperceptible details.
During inference, the process begins with a pure Gaussian noise image $\mathbf{x}_{T}\sim\mathcal{N}(\textbf{{0}},\textbf{{I}})$, the model learns the conditional transition distribution $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{y},\mathbf{a},\mathbf{n})$ and iteratively denoise the image for $T$ steps, generating the target image $\mathbf{x}_{0}$ in the end. Following the notations defined in Sec. III-A, the training and inference processes of our model is summarized in Algorithm 2 and 3.
Algorithm 2 Training our denoising model $\epsilon_{\theta}$
[…]
2:repeat
3:  $(\mathbf{x}_{0},\mathbf{y})\sim q(\mathbf{x}_{0},\mathbf{y})$
4:  $t\sim\text{Uniform}({1,...,T})$
5:  $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
6:  $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$
7:  Take a gradient descent step on
$\nabla_{\theta}\lambda_{t}\left\|\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n}\right)-\epsilon\right\|_{2}^{2}$
8:until converged
Algorithm 3 Inference in $T$ iterative refinement steps
[…]
3:for $t=T,\ldots,1$ do
4:  $z\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $z=0$
5:  $\mathbf{x}_{t-1}=\frac{1}{\sqrt{1-\beta_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t,\mathbf{y},\mathbf{a},\mathbf{n})\right)+\sigma_{t}z$
6:end for
7:return $\mathbf{x}_{0}$
@@ III-D Diffusion-based Restoration Model Framework
IV Experiments
IV-A Dataset
Training Data. Since no publicly available datasets exist for low-light noisy face images, we simulate such data using the CelebAMask-HQ dataset [Lee2020CelebAMaskhq], the InverseISP algorithm [li2024efficient], and the Poisson-Gaussian noise model (Eq. 1).
As shown in Fig. 7, the first step in the simulation process involves passing the RGB image through a pre-trained Inverse ISP network [li2024efficient] to generate a pseudo ground truth RAW mosaic image. For visualization purposes, we display the demosaiced version of this image. We utilize the pretrained weights provided [li2024efficient], which were obtained by training the network on iPhone camera images for unprocessing the RGB images into RAW format.
The pseudo ground truth RAW image then passed through a Poisson-Gaussian simulator to simulate the low-light noisy RAW mosaic image. During the training process, we vary the camera sensor parameters across a wide range of values to allow the model to learn diverse camera sensor noise profiles. The simulation setup assumes a $12$-bit image sensor with parameters similar to typical DSLR cameras. The parameter ranges are summarized in Table V.
TABLE V: Simulation parameters used in the Poisson-Gaussian Simulator.
  Parameter Range
Full Well Capacity (FW) $19000\text{e}^{-}\leq\text{FW}\leq 64000\text{e}^{-}$
Quantum Efficiency ($\mathrm{\eta}$) $0.32\leq\mathrm{\eta}\leq 0.54$
Dark Current ($\mathrm{\sigma_{dark}}$) $2.2\text{e}^{-}\leq\mathrm{\sigma_{dark}}\leq 11.7\text{e}^{-}$
Read Noise ($\mathrm{\sigma_{read}}$) $2.2\text{e}^{-}\leq\mathrm{\sigma_{read}}\leq 10.8\text{e}^{-}$
photo level ($\mathrm{\alpha}$) $13\,\text{ppp}\leq\mathrm{\alpha}\leq 65\text{ppp}$
Testing Data. To test DiffPGD, we utilize both simulated and real-world captured low-light noisy data.
For the simulated testing data, similar to the training data simulation, the simulation setup assumes a $12$-bit image sensor with parameters similar to that of a DSLR camera: a full well capacity (FW) of $36000\text{e}^{-}$, quantum efficiency ($\mathrm{\eta}$) of $0.42$, dark current of $\mathrm{\sigma_{dark}}=7\text{e}^{-}$, and read noise of $\mathrm{\sigma_{read}}=6\text{e}^{-}$. These parameters ensure realistic noise characteristics for evaluating model performance.